OpenStudy (anonymous):

A garden table and a bench cost $450 combined. The cost of the garden table is two times the cost of the bench. What is the cost of the bench?

OpenStudy (anonymous):

Alright, let's list what we have given first ^_^ : - Let's let the garden table be G and bench be B so : - G + B = $ 450 - G = 2 (times) x B Now, RTF (required to find) : - Cost of the bench All you have to do now is substitute G in the general equation I listed above and you'll get :) : G + B = $ 450 2B + B = 450 3B = 450 B = 450/3 B = $ 150 so the cost of the bench = $ 150 ^_^

OpenStudy (anonymous):

2x + x = $450 3x = $450 x = $150

OpenStudy (anonymous):

Let me know if there's something that is unclear to you :)